OpenStudy (rrhtgre):

(D4 - 5D3 + 5D2 + 5D - 6)y = 0 Find the General Solution ( I doubt any of you can haha)

OpenStudy (learner):

`My interpretation:` \(D\) is the linear differential operator. Since no independent variable is given, will call it \(t\). (I'm not going to go through the algebra right now; you should know this part by now ... Yes, it's long!) \(\color{black}{\displaystyle (D^4 - 5D^3 + 5D^2 + 5D - 6)[y] = 0}\) \(\color{black}{\displaystyle (D-1)(D+1)(D+2)(D+3)[y] = 0}\) Next, just a primitive step to write out the general solution. (No repeated roots, no imaginary roots, ... should be pretty easy.)